What's the present value of $1,050 discounted back 5 years if the appropriate interest rate is 6%, compounded monthly?

Select the correct answer.

a. $768.04
b. $762.84
c. $773.24
d. $783.64
e. $778.44
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

1050=P*(1+0.06/12)^(12*5)

P=1050/(1+0.06/12)^(12*5)

=$1050*0.741372196

=$778.44(Approx).